OpenStudy (carolina4567):

That Is the Way the Ball Bounces! Sam completed the following procedure in his Algebra II class: Choose a height from which all of the balls will be dropped one at a time. Vertically along the blank wall, set up the measuring tape and step stool or chair. Have a family member or friend stand on a step stool and drop one of the balls from the chosen height. Drop the ball close enough to the measuring tape to be able to record height, but not touch the tape. Face the measuring tape, opposite the ball's starting point from about 7 or 8 feet high.

OpenStudy (carolina4567):

As the ball falls, measure the height the ball reaches after each bounce for four consecutive bounces. (You may need to repeat the process to ensure that your measurements are accurate. You may choose to video each drop to assure accuracy.) illustration showing a ball bouncing 4 times if someone dropped the ball from 7 or 8 feet high He recorded the height of each bounce, beginning with the height from which the ball originally fell, in the chart below: Ball 1 Description Ball 2 Description Ball 3 Description Height 1 (starting point) 3 ft 3 ft 3 ft Height 2 2.4 1.9 0.9 Height 3 1.9 1.2 0.3 Height 4 1.5 0.8 0.1 Height 5 1.2 0.5 0.02

OpenStudy (carolina4567):

Using complete sentences, answer the following questions: What is the average common ratio between the successive height values of ball 1? Ball 2? Ball 3? Experimental errors may cause common ratios to have some variances within the data for one ball. Use the average common ratio. What is the height of each ball on the fifth bounce (i.e., Height 6)? Use the geometric sequence formula, an = a1rn – 1 and show your work. What is the total distance of the height each ball has traveled in the first five heights? Use the geometric series formula, Sn = the quantity of a sub 1 minus a sub 1 times r to the n power, all over 1 minus r and show your work.

OpenStudy (carolina4567):

Please Help Me !

OpenStudy (carolina4567):

@iGreen @hartnn @AnswerMyQuestions

hartnn (hartnn):

could you start ? where are you stuck?

OpenStudy (carolina4567):

Ok... Um what im not uderstanding is do we create our own data or we use the data given ?

hartnn (hartnn):

the data given, for ball 1 the heights are 3,2.4,1.9,1.5,1.2 find the average common ratio

hartnn (hartnn):

like find the ratio of next height to current height, for all heights and take the average

OpenStudy (carolina4567):

1.25

hartnn (hartnn):

its not 3/2.4 it'll be 2.4/3 =.... 1.9/2.4 =.... 1.5/1.9.... and so on, and then take the average of all the above answers

OpenStudy (carolina4567):

ok so i've got 0.8 , 0.78 , 0.79 , 0.8

OpenStudy (carolina4567):

oops 0.79 , 0.78

hartnn (hartnn):

so what is the average of all those ?

OpenStudy (carolina4567):

0.63

hartnn (hartnn):

can't be ... did you do any calculation mistake?

OpenStudy (carolina4567):

Hmm let me see i added it up and divided by the number of heights

OpenStudy (carolina4567):

3.17 / 5

hartnn (hartnn):

divide it by number of ratios = 4 :)

OpenStudy (carolina4567):

oops sorry

OpenStudy (carolina4567):

0.79

hartnn (hartnn):

thats you common ratio = r for 1st ball do it for ball2 and ball 3 too :)

OpenStudy (carolina4567):

Ok give me a sec

OpenStudy (carolina4567):

Ball 2 is 0.63

hartnn (hartnn):

you want me to verify your answers or are you confident of your workings?

OpenStudy (carolina4567):

I would like you to verify im not usually very good at these things

OpenStudy (carolina4567):

For ball 3 its 1.16

hartnn (hartnn):

go to the next part \(a_n = a_1 r^{n-1}\) a_1 is the first height. r = we just calculated n= 6 (we need 6th height)

hartnn (hartnn):

for each balls, you'll get different answers because a_1 and r are different

OpenStudy (carolina4567):

Ok im going to try

OpenStudy (carolina4567):

Like this right ? an = (3)0.79 ^ 6 -1

hartnn (hartnn):

yes, be careful with the brackets (3)[0.79 ^ (6 -1)] (3)*[0.79 ^ 5]

OpenStudy (carolina4567):

Ohhh Ok thats better

OpenStudy (carolina4567):

0.92

OpenStudy (carolina4567):

0.29 and 6.30

OpenStudy (carolina4567):

Am i correct ?

hartnn (hartnn):

i got 0.954 for first, but 0.92 is close enough

hartnn (hartnn):

0.304 for second, and 0.29 is close enough too :)

hartnn (hartnn):

for ball 3, r is not 1.16 ... try again?

OpenStudy (carolina4567):

ok

OpenStudy (carolina4567):

ok i got it now its 0.29

hartnn (hartnn):

yeah, thats correct so 3*[0.29^5]

OpenStudy (carolina4567):

0.00

hartnn (hartnn):

lol can't be 0.... its 0.006

OpenStudy (carolina4567):

Oh ok :)

OpenStudy (carolina4567):

Sn = 3 - 3 (0.79) 5 / 1 - 0.79 = Sn = 3 - 3 (0.63)5 / 1 - 0.63 = Sn = 3- 3 (0.006)5 / 1 - 0.006

OpenStudy (carolina4567):

Now like ^ that right ?

hartnn (hartnn):

yes, those are correct Sn = [3 - 3 (0.79) 5 ]/ [1 - 0.79] =

OpenStudy (carolina4567):

Is it 0.2

hartnn (hartnn):

[3 - 3 (0.79) 5 ]/ [1 - 0.79] i get 9.889

OpenStudy (carolina4567):

How did you do it ?

hartnn (hartnn):

calculator

OpenStudy (carolina4567):

Mine wont do it ! It gave me 17.42 for the second one

OpenStudy (carolina4567):

Wait 7.30 for the second one

OpenStudy (carolina4567):

And 3.01 for last one

OpenStudy (carolina4567):

THANKKKKK U SO MUCHHHHH!!!! Ur A Life Saver.... Mind Helping Me With One More Question

OpenStudy (carolina4567):

Robbie the Robot is on a weather satellite orbiting Earth about 3600 km above the surface. The Earth’s radius is about 6400 km. He has had a malfunction in his output device, and the satellite is traveling without communication. His last report was only in terms of trigonometric values and was only partially received. It said, sin Θ < 0….. and then he was lost again. Part 1: Create a set of coordinates that would be reasonable for Robbie’s position in space and satisfy his last, partial report. Using complete sentences, describe Robbie’s location and your reasoning.

hartnn (hartnn):

sorry, wish i had that much more time...

OpenStudy (carolina4567):

Oh no worrys you have done enough... Thank you so much !!